# AllSpark: a multimodal spatiotemporal general model

### 31 Dec 2023 v1

@@ 3.4 Modal2Language Bridge
The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality. In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
@@
In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model, and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities. The keys and values in the cross-attention layer are the features outputted by the modal encoders.
@@
The feedforward neural network inherits the classic design from the original Transformer, consisting of two linear layers with an inserted activation layer. The entire process can be formalized as:
$\Phi\left(Q,s_{i}\right)={FFN}\left({softmax}\left(\frac{W_{q}Q^{T}\left(w_{k}s_{i}^{T}\right)^{T}}{\sqrt{D}}\right)*(W_{v}s_{i}^{T})\right)$ (4)
@@
Here, $W_{q}\in R^{D*{hidden}}$, $W_{k}\in R^{d*{hidden}}$, and $W_{v}\in R^{d*{hidden}}$are the linear projection layer weights defined inside the cross-attention layer for $Q$, $K$ and $V$ respectively.
3.5 Task-Guided Text Prompts and Task-Heads
@@
To extend the visual-language multimodal model to 13 spatiotemporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task. Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
@@
We manually designed one to four specific text prompts for each modality. During the training process, to enhance model performance, we employed a strategy of diversifying prompts, randomly selecting one prompt for each forward pass. However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts. Table 2 provides a list of all the text prompts.
Table 2: List of prompts
Dimension
Modal
Prompt
1D
Text
1.Please determine if this movie review is positive or negative?
Code
2D
RGB
1.This remote sensing image belongs to which of the following categories: [Category list of the NWPU dataset] 2.Describe this remote sensing image briefly 3.Find a word that is most relevant to this remote sensing image 4.Describe the key elements in this remote sensing image
MSI
1.Based on the multi-spectral imagery feature description, please classify this object 2.Given the following multi-spectral imagery characteristics, please output the most fitting scene label
HSI
1.Given the spectral information, can you help determine which class this pixel belongs to? 2.Here is the spectral data for a pixel. Considering the typical characteristics of land cover classes, could you provide a detailed analysis and suggest the most likely class for this pixel? 3.The spectral information for a pixel is given, but the data is noisy. Given the potential variability, which land cover classes should be considered as possible candidates for this pixel?
Table
1.Please utilize the provided air quality indicators to accurately predict the concentrations of PM2.5 in the atmosphere 2.Given the air quality indicators, please provide a prediction for the PM2.5 levels at this particular moment
Oblique Photography
1.Please reconstruct the 3D model from the given images of five different views.
Trajectory
1.Based on their past positions and movements in a crowded environment, predict the future trajectory of a selected pedestrian. 2.Using the pedestrian trajectory data, along with additional information about the surrounding environment, predict the future path of the pedestrian. 3.Given the current and past positions of a pedestrian and their neighboring pedestrians, predict the main pedestrian’s trajectory
SAR
1.Based on the SAR imagery feature description, please classify this object 2.Given the following SAR imagery characteristics, please output the most fitting scene label
Infrared
1.Given an infrared image of a person, find and highlight the same person in a set of visible light images, paying close attention to features like clothing, posture, and gait 2.Analyze the characteristics of a person in a visible light image, such as clothing texture, color, and body shape. Then, locate the person with matching features in a series of infrared images
Graph
1.Given the current traffic data including vehicle flow rate, average speed, and time of day from the METRLA dataset, predict the traffic flow 2.Analyze the historical data on vehicle speeds and flow rates from the METRLA dataset for the past week. Identify any patterns or trends and predict the traffic conditions
3D
Video
1.Please analyze this video and describe the characteristics of the action in detail 2.Watch this short clip and provide a step-by-step description of the action sequence
Point Cloud
1.Classify the provided point cloud sample into the correct category 2.Look at the point cloud data characteristics and classify the object 3.Please analyze the given point cloud dataset and determine which category it belongs to. Focus on the shape and structure evident in the point cloud
@@
Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks. To enhance the model’s generalization capability on downstream tasks, we designed specific task heads for each task. Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible. The detailed summary of each modality’s task heads can be found in Table 4.
@@
For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head. For instance, we implemented standard classification tasks on the RGB, MSI, SAR, Video, and PointCloud modalities. Although the task on the HSI modality is a segmentation task, it can be formalized as a per-pixel classification task. Similarly, for the re-identification task on the Infrared modality, it can also be formalized as a classification task. Therefore, a single-layer linear layer is used as task head for the mentioned modalities.
@@
For regression tasks on the Table, Trajectory, and Graph modalities, we also use a linear layer to perform regression predictions. The only difference from classification tasks is the addition of an unscaled operation without learnable parameters. Since the Lynx model itself is a language model, the Code and Text modalities directly use its native text decoder. The task head for the Oblique Photography modality is based on the 3D reconstruction architecture from the Ada-MVSliu2023deep model.
4 Experiment
4.1 Setup
Our experiments aimed to evaluate the adaptability of AllSpark to various spatiotemporal modalities. To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain. The specific details on dataset selection refer to Section 4.2.
Additionally, for result reproducibility, we employed similar experimental settings across all modalities. We used the AdamW optimizer with a learning rate schedule based on cosine annealing. The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments. Table 3 summarizes the hyperparameter settings for experiments on each modality.
Table 3: Summary of hyperparameters
@@ 4.1 Setup
1.0*10-5
3 1
4.2 Dataset
@@
Table 4 summarizes the data and tasks used in experiments for each modality. Below, we will provide detailed explanations in order:
@@
Text: The IMDB dataset is a binary sentiment analysis dataset, consisting of 50,000 reviews from the Internet Movie Database (IMDb) labeled as positive or negative. Additionally, the dataset includes some unlabeled data. In our experiments, only the labeled data from the IMDB dataset was utilized for supervised sentiment classification tasks.
@@
Code: CodeSearchNet is a large-scale dataset of function code and its documentation from GitHub, covering six programming languages: Go, Java, JavaScript, PHP, Python, and Ruby. The task performed on the code modality is Code Document Generation, and we tested it on the Ruby and JavaScript languages.
@@
RGB: NWPU-RESISC45 is a large-scale open dataset for visible light remote sensing image scene classification. It includes 45 land-use categories such as airplane, baseball diamond, beach, commercial area, etc. Each category has 700 remote sensing images, making a total of 31,500 images in the dataset. The image size is 256*256 pixels, and we selected a version of the dataset split by the official release, using 20% of the data for training.
@@
MSI: The EuroSAT dataset is a multispectral modality dataset for land use and land cover (LULC) classification. The samples are sourced from the Sentinel-2 optical satellite and include all 13 bands. The data is categorized into 10 classes, with a total of 27,000 images. We adopted a random 9:1 split for training and testing.
@@
HSI: The Pavia University dataset is a high-spatial-resolution hyperspectral dataset acquired by the ROSIS sensor. It comprises 103 bands with a size of 610*340 pixels. The dataset includes nine land cover categories such as Asphalt, Meadows, Gravel, etc. We used a 4:6 split for training and testing.
@@
Table: The PRSA dataset is a collection of air quality data from multiple stations in Beijing, containing hourly measurements of air pollutants. The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations. In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, wind direction). The task was to predict the concentration of PM2.5. We split the data into training (40%) and testing sets.
@@
Trajectory: The ETH-UCY dataset is one of the widely used benchmarks for pedestrian trajectory prediction, divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2. In our experiments, we utilized the ETH subset.
@@
SAR: The MSTAR dataset is a synthetic aperture radar dataset designed for military stationary target recognition, comprising ten categories of military targets. We employed the standard operating conditions (SOC) dataset preprocessing method proposed by S. Chen et al. chen2016target , ensuring that the serial numbers and target configurations are consistent between the test and training sets, while the aspects and depression angles differ.
@@
Infrared: SYSU-MM01 is a cross-modal human re-identification dataset with optical-infrared modalities, including images from 4 RGB cameras and 2 infrared cameras of 491 different individuals. We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals, while the test set consists of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for re-identification targets.
@@
Graph: METR-LA is a traffic data collection from loop detectors on the highways of Los Angeles, spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
@@
Oblique Photography: WHU-OMVS is an oblique photography dataset designed for 3D reconstruction tasks. The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information. It consists of six areas, and in our experiments, we use area 1 as the training set and area 2 as the test set.
@@
Video: UCF101 dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips. The videos have a combined duration of 27 hours and a resolution of 320*240 pixels, sourced from YouTube.
@@
PointCloud: ModelNet40 is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects. We follow the official dataset split, with 9,843 objects used for training and 2,468 for testing.
4.3 Results
4.3.1 RGB
We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the Top-1 Accuracy as the evaluation metric. AllSpark leverages expert knowledge from Lynx by loading its pretrained weights. Therefore in Table 4, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset. The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to the SOTA (95.69). This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
Table 4: RGB scene classification with AllSpark(80% as trainset)
@@ 4.3.1 RGB
SeCo-ResNet-50manas2021seasonal
@@
4.3.2 MSI
We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset. In the experiment, all 13 spectral bands of the images were simultaneously input into the model. The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was Top-1 Accuracy. AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups in Table 5: those with expert knowledge intervention and those without. Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset. The results show that our model outperforms most models in the no expert knowledge group, with a margin of only 2.60 compared to the state-of-the-art model (ResNet-152). Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
Table 5: MSI Land Cover Classification with AllSpark
Expert Knowledge
@@ 4.3.2 MSI
4.3.3 HSI
We conducted pixel classification task on the Pavia University dataset for the hyperspectral modality. The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel. The reported metrics include Overall Accuracy (OA), Average Accuracy (AA), and Kappa. Since AllSpark does not possess expert knowledge in the hyperspectral modality, we compared it with the semi-supervised baselines summarized by D. Uchaev and D. Uchaev uchaev2023small . The results in Table 6 demonstrate that AllSpark outperforms many hyperspectral image classification methods such as IFRF and S-DMM, with a difference of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
Table 6: HSI pixel Classification with AllSpark
@@ 4.3.3 HSI
DCFSL
90.71 90.2 87.73
TC-GAN
93.2 91.6 91
PRNet-RF
95.6 94.96 94.27
4.3.4 Table
Table 7: PM2.5 prediction with AllSpark
Expert Architecture Method PM2.5

### 25 Feb 2024 v2

@@ III-D Modal2Language Bridge
Although the modal encoders have transformed data from various heterogeneous modalities into a unified token sequence, there are still differences in dimensions between different modal tokens, making it difficult to perceive by a multimodal large language model. The modal bridge, based on the Lynx model, aims to perform dimensional projection from tokens of various modalities to tokens of the language modality. In its implementation, the modal bridge consists of stacked layers, including cross-attention layers and feedforward neural network layers.
@@
In the cross-attention layer, we predefine a learnable query vector $Q\in R^{N*D}$, where $D$ is the internal dimensionality of the language model and $N$ serves as a hyperparameter that can be flexibly adjusted to accommodate inputs from different modalities. The keys and values in the cross-attention layer are the features outputted by the modal encoders.
@@
The feedforward neural network inherits the classic design from the original transformer and consists of two linear layers with an inserted activation layer. The entire process can be formalized as follows:
$\Phi\left(Q,s_{i}\right)={FFN}\left(\sigma\left(\frac{QW_{q}^{T}\left(s_{i}W_{k}^{T}\right)^{T}}{\sqrt{D}}\right)s_{i}W_{v}^{T}\right)$ (4)
@@
Here, $W_{q}\in R^{D*{hidden}}$, $W_{k}\in R^{d*{hidden}}$, and $W_{v}\in R^{d*{hidden}}$ are the linear projection layer weights defined inside the cross-attention layer for $Q$, $K$ and $V$, respectively. $\sigma$ denotes the softmax operation.
III-E Task-Guided Text Prompts and Task Heads
@@
To extend the visual-language multimodal model to thirteen spatio-temporal modalities without intervention from modality expert knowledge, we designed specific text prompts and task heads for each modality and task. Text prompts are used to guide the multimodal language model in correctly interpreting each modality’s data, while task heads are employed to match the model’s parsing results with specific downstream tasks.
@@
We manually designed one to four specific text prompts for each modality. During the training process, to enhance model performance, we employed a strategy of diversifying prompts, randomly selecting one prompt for each forward pass. However, during testing, for the sake of result stability and reproducibility, the prompt was fixed to be the first prompt among all prompts. Table XVI provides a list of all the text prompts.
@@
Additionally, we noticed that there is still a gap between the multimodal language model’s parsing results and the specific downstream tasks. To enhance the model’s generalization capability on downstream tasks, we designed specific task heads for each task. Moreover, to ensure the transferability of modalities across different tasks, the design principle for task heads is to be as simple and lightweight as possible. A detailed summary of each modality’s task heads can be found in Table 4.
@@
For classification tasks or downstream tasks that can be formalized as classification tasks, we uniformly use a simple single-layer linear layer as the task head. For instance, we implemented standard classification tasks on the RGB, MSI, SAR, video, and PointCloud modalities. Although the task involving the HSI modality is segmentation, it can be formalized as a per-pixel classification task. Similarly, the reidentification task in the infrared modality can also be formalized as a classification task. Therefore, a single-layer linear layer is used as the task head for the mentioned modalities.
@@
For regression tasks on the table, trajectory, and graph modalities, we also use a linear layer to perform regression predictions. The only difference from classification tasks is the addition of an unscaled operation without learnable parameters. Since the Lynx model itself is a language model, the code and text modalities directly use its native text decoder. The task head for the oblique photography modality is based on the 3D reconstruction architecture from the Ada-MVS[42] model.
IV Experiment
IV-A Setup
TABLE II: Summary of hyperparameters
@@ IV-A Setup
$1.0\times 10^{5}$
3
1
Our experiments evaluated the adaptability of AllSpark to various spatio-temporal modalities. To provide a realistic evaluation of its multimodal adaptability and facilitate comparisons with existing work in the field, we followed the principles of simplicity, reproducibility, and choosing datasets and tasks that have been widely studied in the domain. The specific details on dataset selection can be found in Section IV-B.
Additionally, for reproducibility, we employed similar experimental settings across all modalities. We used the AdamW optimizer with a learning rate schedule based on cosine annealing. The hyperparameters were adjusted slightly in terms of training epochs and learning rates for different experiments. Table II summarizes the hyperparameter settings for the experiments on each modality.
IV-B Dataset
@@
Table 4 summarizes the data and tasks used in the experiments for each modality. Below, we will provide detailed explanations in order:
@@
Text: The IMDB[43] dataset is a binary sentiment analysis dataset consisting of 50,000 reviews from the Internet Movie Database (IMDb) labelled as positive or negative. Additionally, the dataset includes some unlabelled data. In our experiments, only the labelled data from the IMDB dataset were utilized for supervised sentiment classification tasks.
@@
Code: CodeSearchNet[44] is a large-scale dataset of function code and its documentation from GitHub that covers six programming languages: Go, Java, JavaScript, PHP, Python, and Ruby. The task performed on the code modality is code document generation, and we tested it on the Ruby and JavaScript.
@@
RGB: NWPU-RESISC45[45] is a large-scale open dataset for visible light remote sensing image scene classification. The dataset included 45 land use categories, such as airplanes, baseball diamonds, beaches, and commercial areas. Each category included 700 remote sensing images, for a total of 31,500 images. The image size is 256*256 pixels, and we selected a version of the dataset split by the official release using 20% of the data for training.
@@
MSI: The EuroSAT[46] dataset is a multispectral dataset for land use and land cover (LULC) classification. The samples are sourced from the Sentinel-2 optical satellite and include all 13 bands. The data are categorized into 10 classes for a total of 27,000 images. We adopted a random 9:1 split for training and testing.
@@
HSI: The Pavia University dataset is a high-spatial-resolution hyperspectral dataset acquired by the ROSIS sensor. It comprises 103 bands with a size of 610*340 pixels. The dataset includes nine land cover categories, such as asphalt, meadows, and gravel. We used a 4:6 split for training and testing.
@@
Table: The PRSA[47] dataset is a collection of air quality data from multiple stations in Beijing that contains hourly measurements of air pollutants. The data spans from March 1, 2013, to February 28, 2017, across 12 monitoring stations. In our experiments, we used various features, including time, station information, 4 air pollutant variables (SO2, NO2, CO, and O3), and 6 meteorological variables (temperature, pressure, dew point temperature, amount of precipitation, wind speed, and wind direction). The task was to predict the concentration of PM2.5. We split the data into training (40%) and testing sets.
@@
Trajectory: The ETH-UCY[48, 49] dataset is a widely used benchmark for pedestrian trajectory prediction and is divided into five subsets: ETH, HOTEL, UNIV, ZARA1, and ZARA2. In our experiments, we utilized the ETH subset.
@@
SAR: The MSTAR[50] dataset is a synthetic aperture radar dataset designed for military stationary target recognition that comprises ten categories of military targets. We employed the standard operating conditions (SOCs) dataset preprocessing method proposed by S. Chen et al. [22], ensuring that the serial numbers and target configurations are consistent between the test and training sets while the aspects and depression angles differ.
@@
Infrared: SYSU-MM01[51] is a cross-modal human reidentification dataset with optical-infrared modalities that includes images from 4 RGB cameras and 2 infrared cameras of 491 different individuals. We adopted the official dataset split, with the training set containing 20,284 RGB images and 9,929 infrared images from 296 individuals and the test set consisting of 3,803 infrared images from 96 individuals used as queries and 301 randomly selected RGB images for reidentification targets.
@@
Graph: METR-LA is a traffic dataset collected from loop detectors on the Los Angeles highways spanning from March 1, 2012, to June 30, 2012. The task is traffic flow prediction.
@@
Oblique Photography: WHU-OMVS[42] is an oblique photography dataset designed for 3D reconstruction tasks. The dataset provides imagery from five different viewpoints, along with camera parameters and other relevant information. It consists of six areas, and in our experiments, area 1 is used as the training set, and area 2 is used as the test set.
@@
Video: The UCF101[52] dataset is a human action recognition dataset comprising 101 action classes with a total of 13,320 video clips. The videos have a combined duration of 27 hours and a resolution of 320*240 pixels and were sourced from YouTube.
@@
PointCloud: ModelNet40[53] is a synthetic point cloud dataset consisting of 40 object categories and a total of 12,311 point cloud objects. We follow the official dataset split, with 9,843 objects used for training and 2,468 for testing.
IV-C Results
IV-C1 RGB
TABLE III: rgb image classification with allspark
@@ IV-C1 RGB
SeCo-ResNet-50[58]
@@
We evaluated the performance of AllSpark on the RGB image scene classification task using the NWPU-RESISC45 dataset, with the top-1 accuracy as the evaluation metric. AllSpark leverages expert knowledge from Lynx by loading its pretrained weights. Therefore, in Table III, we compare AllSpark with the state-of-the-art models known on the NWPU-RESISC45 dataset. The results indicate that AllSpark outperforms most baseline models, with a margin of only 0.84 compared to that of the SOTA (95.69). This highlights AllSpark’s exceptional perception and interpretation capabilities in the RGB modality.
IV-C2 MSI
We evaluated the performance of AllSpark on the MSI scene classification task using the EuroSAT dataset. In the experiment, all 13 spectral bands of the images were simultaneously input into the model. The model’s objective was to correctly classify the images into one of the 10 specified categories, and the evaluation metric chosen was the top-1 accuracy. AllSpark does not possess expert knowledge in the multispectral modality, so we categorize the baseline models into two groups, as shown in Table IV: those with expert knowledge intervention and those without. Expert knowledge intervention refers to baseline models pretraining on large datasets such as BigEarthNet and then fine-tuning on the EuroSAT dataset, while no expert knowledge indicates baseline models trained directly from scratch on the EuroSAT dataset. The results show that our model outperforms most models in the no expert knowledge group, with a margin of only 2.60 compared to the state-of-the-art model (ResNet-152). Furthermore, AllSpark lags behind the best result in the expert knowledge group by only 4.75, demonstrating its excellent adaptability to the multispectral modality.
TABLE IV: msi land cover classification with allspark
Expert Know.
@@ IV-C2 MSI
IV-C3 HSI
We conducted a pixel classification task on the Pavia University dataset for the hyperspectral modality. The model treats all spectral bands of a single pixel as one sample and predicts the land cover category of that pixel. The reported metrics include overall accuracy (OA), average accuracy (AA), and kappa. Since AllSpark does not possess expert knowledge of the hyperspectral modality, we compared it with the semisupervised baselines summarized by D. Uchaev and D. Uchaev [72]. The results in Table V demonstrate that AllSpark outperforms many hyperspectral image classification methods, such as IFRF and S-DMM, by a factor of 6.42 compared to the best result in terms of OA, highlighting AllSpark’s superior adaptability to the hyperspectral modality.
TABLE V: hsi pixel classification with allspark
@@ IV-C3 HSI
DCFSL 90.71 90.2 87.73
TC-GAN 93.2 91.6 91
PRNet-RF 95.6 94.96 94.27
IV-C4 Table
For the table modality, we evaluated AllSpark on the regression prediction task using the PRSA[47] dataset. The task involves predicting the concentration of PM2.5 in the air using features such as time, site, 4 air pollutants, and 6 meteorological variables (as detailed in Section IV-B). The performance metrics include the root mean squared error (RMSE), mean absolute error (MAE), and R-squared (R2).
TABLE VI: PM2.5 prediction with allspark
Expert Arch. Method PM2.5